OpenStudy (anonymous):

the width of a rectangle playground is 18 meters less then the length find the width.when the perimeter of the rectangle is 84 meters

OpenStudy (anonymous):

You would set up the equation as (x)(x-18)=84.

OpenStudy (raden):

first, solve for x of expression bellow! 2(x+x-18) = 84 i assumed the length is x and then the width be (x-18)

OpenStudy (anonymous):

so i got that but im lost now

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

im so cofused do you mind helping please

jhonyy9 (jhonyy9):

let the length being l so than the width will be l-18 but you know that the perimeter of a rectangle you get assuming the length with width two time so wath mean that P = 2l + 2w but you know that w=l-18 so than P=2l+2w=2l+2(l-18)=2l+2l-2*18=4l-36=4(l-9)

OpenStudy (anonymous):

thank you so much

jhonyy9 (jhonyy9):

yw was my pleasure good luck bye